OpenStudy (anonymous):

five times a number k minus 9 is two-third of a number j. solve for j

OpenStudy (amistre64):

five times [a number k minus 9] "is" [two-third] of a number j

OpenStudy (amistre64):

since there is not comma between k and minus, it should be regarded as a single unit

OpenStudy (anonymous):

what is it no 6k-93/2=j?

OpenStudy (amistre64):

five times [a number k minus 9] "is" [two-third] of a number j 5 * [k-9] = 2/3 * j 5(k-9)= (2/3)j 5(3)(k-9)/2= j